# Aligning Machine and Human Visual Representations across Abstraction Levels

### v3

@@ Abstract
Deep neural networks have achieved success across a wide range of applications, including as models of human behavior in vision tasks. However, neural network training and human learning differ in fundamental ways, and neural networks often fail to generalize as robustly as humans do, raising questions regarding the similarity of their underlying representations. What is missing for modern learning systems to exhibit more human-like behavior? We highlight a key misalignment between vision models and humans: whereas human conceptual knowledge is hierarchically organized from fine- to coarse-scale distinctions, model representations do not accurately capture all these levels of abstraction. To address this misalignment, we first train a teacher model to imitate human judgments, then transfer human-like structure from its representations to refine the representations of pretrained state-of-the-art vision foundation models via finetuning. These human-aligned models more accurately approximate human behavior and uncertainty across a wide range of similarity tasks, including a new dataset of human judgments spanning multiple levels of semantic abstractions. They also perform better on a diverse set of machine learning tasks, increasing generalization and out-of-distribution robustness. Thus, infusing neural networks with additional human knowledge yields a best-of-both-worlds representation that is both more consistent with human cognition and more practically useful, thus paving the way toward more robust, interpretable, and human-like artificial intelligence systems.
@@
AI alignment, human cognition, representation learning, computer vision
@@ 1 Introduction
Figure 1: a: An example of the triplet odd-one-out task where a human and a neural network model choose the same (top) and a different (bottom) odd-one-out image respectively. b: The different parts of the AligNet framework depicted from end to end. First, we develop a teacher model of human judgments using the THINGS dataset (first panel). Second, we apply this model to ImageNet and cluster its latent representations into semantically meaningful categories (second panel). This allows us to generate arbitrarily many similarity judgments. Third, the obtained human-like similarity structure information is distilled into a student vision foundation model using a novel loss function (third panel). c: Representative human alignment (top panel) and machine learning downstream (bottom panel) results show significant performance improvements of the aligned over the non-aligned version of the ViT-B classifier (up to 123% of relative improvement for ML downstream transfer). d: 2D latent space projection for visualizing the change in the representations after alignment. While the representations of a standard ViT-B classifier model are unstructured and categories overlap, after alignment the representations are grouped into meaningful categories.
@@
While deep learning has recently driven rapid progress in areas of artificial intelligence such as natural language processing [LeCun et al., 2015; Brown et al., 2020] and computer vision [Radford et al., 2021; Dosovitskiy et al., 2021; Zhai et al., 2023; Kirillov et al., 2023], even the best of these systems often fail in ways that humans would not [Lapuschkin et al., 2016; Geirhos et al., 2018; Lapuschkin et al., 2019; Geirhos et al., 2020; Hermann et al., 2020; Thrush et al., 2022; Bowers et al., 2022; Muttenthaler et al., 2023b; Hermann et al., 2023]. These failures have led to renewals [Lake et al., 2017; Bowers et al., 2022] of older arguments that neural networks lack essential ingredients of human intelligence [Fodor and Pylyshyn, 1988; Marcus, 1998; McClelland et al., 2010; Holyoak and Hummel, 2014; Tenenbaum et al., 2011]. How can we build systems that have more human-like behavior?
@@
Human perception is robust to changes in the environment and generalizes across different visual settings [Lake et al., 2015, 2017; Roads and Mozer, 2017; Geirhos et al., 2018; Lake and Baroni, 2023]. However, model performance declines—often drastically—if the data distribution between training and test sets is shifted [e.g., Sugiyama et al., 2007; Sugiyama and Kawanabe, 2012; Geirhos et al., 2020; Hendrycks et al., 2021a]. This lack of robustness in the representations of vision models poses a challenge for many downstream applications that require the ability to generalize [e.g., Lapuschkin et al., 2019; Pooch et al., 2020; Geirhos et al., 2020]. In addition, humans tend to be well-calibrated—for example, when they are asked to judge visual similarity [Roads and Mozer, 2017]—that is, humans’ (un)certainty tends to correlate with their (in)accuracy. AI systems, on the other hand, are often overconfident and show high certainty [Guo et al., 2017] even when their predictions are incorrect [Minderer et al., 2021]. Thus, many differences remain to be reconciled before we can ultimately achieve human-like artificial intelligence.
@@
In this work, we highlight a key misalignment between humans and deep learning models that may underlie some of these differences: model representations tend to fail to capture the multi-level conceptual structure of human perception. Although model representations successfully encode the human-perceived similarity structure among entities that share local visual and semantic features (e.g., textures or colors of different breeds of dog), the global relationships between concepts that are more visually and semantically distinct (e.g., dogs and fish, which are both animate but are not visually similar at all) are modeled much less systematically. Human neural representations, however, are organized by global features like animacy [e.g., Kriegeskorte et al., 2008b; Cichy et al., 2019; Hebart et al., 2020], and at multiple finer scales that capture nuanced semantic relationships [Connolly et al., 2012; Cichy et al., 2019; King et al., 2019]. Could this lack of global organization in the representations of deep learning models across levels of the conceptual hierarchy contribute to the aforementioned weaknesses of these models?
One key step to assessing—and addressing—this misalignment is to be able to systematically and reliably compare human and machine judgments, e.g., of similarity or categorical membership. Such comparisons are daunting, however, because of the substantial difference in data availability between the two sources. Human behavioral data is scarce because collecting representative samples is challenging and expensive [Cichy et al., 2019; King et al., 2019; Hebart et al., 2020, 2023]. The cost of getting (human) similarity judgments scales with the square of the number of images. For this reason, the largest human similarity-judgment datasets to date are several orders of magnitudes smaller than the datasets used to pretrain deep learning models. To build more accurate models of human perception, we propose a novel framework that enables the simulation of arbitrarily large (albeit imperfect) human-like similarity judgment datasets for alignment via a teacher model. We use the created datasets to fine-tune large pretrained neural network models to yield more human-like representations and behavior. This approach allows substantially improving human-model alignment of a broad range of vision foundation models, without repeating the entire pretraining process.
We begin by using an affine transformation to align neural network model representations with human semantic judgments in triplet odd-one-out tasks [Muttenthaler et al., 2023a]. In doing so, we regularize the alignment to preserve the local representation structure of the models, and additionally leverage uncertainty measures over human responses to improve model calibration. This enables us to define a transformation toward more human-like representations of a state-of-the-art image/text vision foundation model (VFM). For the remainder of this paper, we use this aligned version of the VFM as a surrogate model for generating human similarity judgments. This approach is loosely analogous to the use of surrogate models for pseudo-labelling in semi-supervised classification learning [e.g. Lee et al., 2013]. With our surrogate model, we hypothesize that we can create a larger training dataset with multiple levels of conceptual granularity that will serve to provide cross-level constraints on human alignment.
To do so, we group the representations of the surrogate model into meaningful superordinate categories which we use to sample semantically meaningful triplets. Subsequently, we obtain odd-one-out responses from our surrogate model for all the triplets that we generate. This results in a large dataset of human-like triplet odd-one-out responses. We call this dataset AligNet. To transfer this carefully curated human-like similarity structure into a pretrained neural network foundation model, we introduce a new Kullback-Leibler divergence based objective function that facilitates the learning process. The different steps of the AligNet framework are visualized in Fig. 1c. While we use the ImageNet dataset [Russakovsky et al., 2015] because it is publicly available and widely used in the cognitive science and computer vision communities, our framework allows for different versions of AligNet to be created based on other image datasets, irrespective of whether the datasets are labelled.
To validate that AligNet can indeed help to increase the alignment between models and humans, we used crowd-sourcing to collect a novel evaluation dataset of human semantic judgments across multiple levels of abstraction that we call Levels. As anticipated, we find that models are worst at predicting human judgments that involve relationships among entities that have little visual or semantic overlap and are thus considered more abstract. We address this limitation by making models more human-aligned and correspondingly more robust. We demonstrate that the global or coarse-grained representational alignment instilled via AligNet—which we refer to as soft-alignment—improves model alignment with human judgments across a variety of visual similarity tasks. Additionally, we find that model uncertainty is better calibrated to human uncertainty. Moreover, soft-alignment substantially improves accuracy and robustness across many downstream machine learning tasks (see Fig. 1c), thus increasing the generalization capabilities of the model representations.
In addition, AligNet yields a suite of open-source adaptations of vision foundation models that encode hierarchical human knowledge through soft-alignment [^1] Models will be open-sourced upon acceptance of the manuscript for publication.. Its general significance stems from enabling the sciences, medicine, and industries to use more human-like and robust vision models for any downstream application. The use of the ImageNet data to create AligNet gives a publicly available image dataset that can be employed by any researcher or practitioner without restriction.
@@
In summary, our work contributes to better understanding the key differences between artificial and natural intelligence. Moreover, our results illustrate a principle for aligning models and humans—focusing on the multi-resolution relational structure of human knowledge—that may be critical to tackling the more general problem of achieving human-like artificial intelligence.
@@ 2 Results
What is missing in modern learning systems to exhibit more human-like behavior? We approach this question jointly from the perspectives of cognitive science and artificial intelligence. Specifically, our results below establish (1) a mathematical framework for aligning machine learning representations with human similarity judgments (see Sec. 2.1 and Sec. 4.1); and demonstrate that (2) our methodological framework, soft-alignment, indeed gives rise to systems with more human-like behavior in object similarity tasks by teaching models about the hierarchical structure of human conceptual knowledge (see Sec. 2.2). We demonstrate this using (i) human object similarity experiments using triplet odd-one-out or multi-arrangement tasks (see Sec. 2.2.1) and (ii) a new, carefully curated dataset, Levels, that allows us to systematically study the differences between human and deep learning model similarity spaces across three levels of knowledge abstraction (see Sec. 2.2.2). Moreover, we find that (3) soft-alignment leads to models that better reflect the hierarchical structure of human conceptual knowledge (see Sec. 2.3) and show that (4) soft-alignment improves generalization and robustness of models for a wide range of computer vision tasks (see Sec. 2.4).
2.1 Machine learning for alignment
To build models with more human-like behavior it is necessary to inject additional supervision about human behavior into the representations of foundation models. One way to achieve this is via an objective function that explicitly forces a model’s representation to inherit a human concept space [Muttenthaler et al., 2023a, b; Fu et al., 2023]. Previous work has shown that the objective function in addition to the (pre-)training data is the most central variable in the pursuit of human-like vision models [Muttenthaler et al., 2023a; Conwell et al., 2022].
We used the THINGS dataset [Hebart et al., 2020] — a behavioral dataset of human responses in a triplet odd-one-out task for object images — to build a surrogate model (denoted as teacher) that approximates a human object similarity space. This similarity space is determined by the human odd-one-out responses from which abstract human concepts can be inferred [Hebart et al., 2020; Muttenthaler et al., 2022]. To better match the human responses, we learned an affine transformation of the pretrained teacher model’s representation space of the form: ${\bm{x}}^{\prime}={\bm{W}}{\bm{x}}+{\bm{b}}$, where ${\bm{x}}\in\mathbb{R}^{p}$ is the model’s latent representation for an object image in the data, and ${\bm{W}}$ and ${\bm{b}}$ are a learned transformation matrix and a bias vector respectively. The pairwise similarities of the teacher model between two object images $i,j$ are modeled as the dot product between their corresponding transformed representations: $S^{\prime}_{i,j}\coloneqq\left({\bm{W}}{\bm{x}}_{i}+{\bm{b}}\right)^{\top}\left({\bm{W}}{\bm{x}}_{j}+{\bm{b}}\right)$. We obtained the affine transformation by maximizing the probability to select a human choice $c\coloneqq\{a,b\}$ given a triplet of object indices $t\coloneqq\{i,j,k\}$, where $c\subset t$, and the corresponding object representations of the pretrained teacher model over $n$ images:
$\operatorname*{arg\,min}_{{\bm{W}},{\bm{b}}}-\frac{1}{n}\sum_{s=1}^{n}\log p\left(c_{s}|t_{s},{\bm{S}}^{\prime}\right)+\Gamma({\bm{W}}),$
where $\Gamma({\bm{W}})$ performs $\ell_{2}$-regularization of the transformation matrix ${\bm{W}}$ toward the scaled identity matrix to preserve the local similarity structure of the original representation space (see Eq. 3 for more details about the type of regularization). We used stochastic gradient descent (SGD) to find the transformation variables ${\bm{W}}$ and ${\bm{b}}$. See Sec. 4.1 and the Supplementary Material for more details about the optimization and how the affine transformation is learned.
After we found a transformation that best matches the human responses, we clustered the transformed teacher representations into meaningful superordinate categories (see Sec. 4.1 and the Supplementary Material). We used these clusters to sample triplets—two objects from one cluster and one from a different cluster—for generating a synthetic dataset of human-like triplet odd-one-out responses. To transfer the structure of the pairwise similarity matrix of the teacher, ${\bm{S}}^{\prime}$, into a different student neural network $f_{{\theta}^{\dagger}}$ (e.g., a Transformer-based VFM), we introduced a novel Kullback-Leibler divergence based objective function similar to Eq. 2 that facilitates the distillation process. This loss function is defined as
$\mathcal{L}_{\mathrm{alignet}}({\bm{S}}^{\prime},{\bm{S}}^{\dagger})\coloneqq\frac{1}{B}\sum_{s=1}^{B}\log\sigma\left(\left[S^{\prime}_{i,j},S^{\prime}_{i,j},S^{\prime}_{j,k}\right],\tau^{\prime}\right)_{s}-\sigma\left(\left[S^{\prime}_{i,j},S^{\prime}_{i,j},S^{\prime}_{j,k}\right],\tau^{\prime}\right)_{s}\log\sigma\left(\left[S^{\dagger}_{i,j},S^{\dagger}_{i,k},S^{\dagger}_{j,k}\right],\tau^{\dagger}\right)_{s},$
where ${\bm{S}}^{\prime}$ and ${\bm{S}}^{\dagger}$ are the pairwise similarity matrices of the teacher and the student representation spaces respectively, $\sigma$ is a $\mathrm{softmax}$ function that transforms similarities into probabilities with temperature values $\tau^{\prime}$ and $\tau^{\dagger}$ for the teacher and student respectively, and $B$ is the total number of samples used for the similarity transfer. The final AligNet objective is defined as
where $\theta^{*}$ are the parameters of the pretrained student before fine-tuning by alignment and $\theta^{\dagger}$ are the parameters of the fine-tuned student. The goal of the right-hand side of the equation is to preserve the fine-grained structure of the pretrained representation space as a function of $\lambda$ which determines the strength of the regularization. See Fig. 1b for a visualization of the different steps in the AligNet framework. In the remainder of this manuscript, we refer to the approach of providing supervision regarding human object similarity via a surrogate teacher model as soft-alignment.
2.2 Toward human-like systems via a surrogate approach
@@
In the following, we demonstrate that soft-alignment gives rise to systems with more human-like behavior. We evaluated the alignment between human mental and neural network representations across a wide range of object similarity tasks using representational similarity analysis [RSA; Kriegeskorte et al., 2008a] and by comparing responses in a triplet odd-one-out task [Muttenthaler et al., 2023a, b; Fu et al., 2023]. As in previous work [Cichy et al., 2019; King et al., 2019; Hebart et al., 2020; Muttenthaler et al., 2023a, b], we used the Spearman rank correlation coefficient to quantify the similarity of human and model representational similarity matrices (RSMs; for details see Sec. 4.2 or the Supplementary Material). As a measure of accuracy for the triplet odd-one-out task, we defined the degree of alignment as the fraction of triplets for which models selected the same odd-one-out choice as the human participants.
First, we validated that our surrogate model performs close to the human noise-ceiling of $66.67\%$ for the THINGS triplet odd-one-out data [Hebart et al., 2020, 2023]. After applying the Uncertainty Distillation optimization (see Eq. 3) to the teacher model representations (see A.1.4 in the Supplementary Material), we observed a quantitative agreement with the human odd-one-out responses in $61.7\%$ of all triplets in the THINGS data (see second dashed line in Fig. 2a). Second, given the triplet odd-one-out responses generated with our surrogate model, we found that after fine-tuning vision foundation models (VFMs) on that data (see Sec. 2.1), all models showed significantly improved alignment with human judgments, irrespective of the pretraining task and the objective function used for training the base model, the cognitive similarity task, and the similarity metric (see Fig. 2a-c).
Notably, we observed relative increases in performance between $21.28-74.47\%$ across a number of representative VFMs trained on different tasks—image captioning, classification, self-supervised learning, image/text alignment—with different training objectives—supervised, unsupervised, self-supervised. Our findings generalize across various object similarity tasks that are commonly used in the cognitive sciences.
Triplet odd-one-out task. Models fine-tuned on AligNet achieved odd-one-out accuracies of $60.45\%$ (ViT-B) -- $62.54\%$ (SigLIP ViT-B) of which $62.54\%$ is the best alignment score of a neural network model on the THINGS dataset to date [cf. Muttenthaler et al., 2023a, b]. In contrast, the base model representations achieved odd-one-out accuracies of $34.87\%$ (ViT-B) -- $51.21\%$ (DINOv2 ViT-B). The odd-one-out accuracy of ViT-B increased substantially from $34.87\%$ to $60.45\%$. This is a relative increase in performance of $73.35\%$. All performance increases are statistically significant at $\alpha=0.05$.
It is not surprising that AligNet fine-tuning significantly improved the alignment of models with the THINGS triplet odd-one-out choices because the affine transformation that is part of the distillation process was learned on that data (see Sec. 2.1 and Sec. 4.1). Thus, we validated whether models fine-tuned on AligNet generalize to other human similarity judgment datasets beyond the THINGS triplet dataset.
Other object similarity tasks. We performed our evaluation on the same human similarity judgment datasets that were analyzed in Muttenthaler et al. [2023b]. The human similarity judgments were collected by either asking participants to arrange natural object images from different categories on a computer screen [Cichy et al., 2019; King et al., 2019] or rate pairwise similarities on an ordinal Likert scale [Peterson et al., 2018]. See the Supplementary Material for more details about the datasets and how they were collected.
We found that AligNet fine-tuned models are substantially better-aligned with human similarity judgments than their base models and even better than an affine transformation learned on the THINGS dataset (see the Supplementary Material for the latter finding). This observation holds irrespective of the similarity task—multi-arrangement [Cichy et al., 2019; King et al., 2019] or Likert-scale pairwise similarity ratings [Peterson et al., 2016, 2018]—that was used to infer object similarity from the human subjects. See Fig. 2b,c for how much AligNet fine-tuning improved the degree of alignment upon their base model representations, measured as Spearman rank correlation between the upper triangulars of model and human RSMs (see Sec. 4.2 for how RSMs are obtained). While the base model representations achieve Spearman rank correlation coefficients of $\rho=0.045-0.508$ and $\rho=0.223-0.477$ with the human responses for coarse-grained and fine-grained object similarity respectively, AligNet models showed a statistically significant improvement in the correlation coefficients and achieved Spearman rank correlations of $\rho=0.655-0.760,p<0.001$ for coarse-grained and $\rho=0.499-0.561,p<0.001$ for fine-grained object similarity, depending on the dataset (Peterson et al. [2016, 2018]: $\rho=0.670-0.717,p<0.001$ for coarse, $\rho=0.499-0.561,p<0.001$ for fine; Cichy et al. [2019]: $\rho=0.747-0.760,p<0.001$; King et al. [2019]: $\rho=0.655-0.666,p<0.001$. The improvements for the classifier are particularly striking and range from a $6.3$ fold increase in Spearman rank correlation coefficient for Peterson et al. [2016, 2018] to a $14.47$ fold increase for King et al. [2019].
2.2.1 The Levels dataset of human similarity judgments
To systematically validate that soft-alignment yields model representations that reflect the hierarchical structure of human conceptual knowledge, we collected a new multi-resolution similarity judgment dataset called Levels. This dataset was based on the triplet odd-one-out task and included judgments on three different types of triplets: global coarse-grained semantic, which require deciding on the odd-one-out in broadly different categories; local fine-grained semantic, involving discerning subtle distinctions within the same category; and class-boundary, testing the capacity to identify category boundaries. For details about Levels see Sec. 4.3.
Descriptive statistics for human response times and uncertainty measures. We examined the human response time (RT) differences across the three abstraction settings as an indicator of the cognitive processing demands required to select the odd-one-out (see Fig. 5A). As expected, responses were fastest in the class-boundary triplets (Mdn = 2.05 s, SE = 0.01), which probed decisions with relatively low ambiguity. RTs for fine-grained semantic triplets were significantly longer (Mdn = 3.01 s, SE = 0.01; $t(447)=48.784$, $p<0.001$), and the longest RTs were observed for the coarse-grained semantic triplets (Mdn = 3.25 s, SE = 0.01; $t(447)=15.036$, $p<0.001$ compared to fine-grained). An explanation for the longer RTs in the latter conditions can be that these triplets involved complex, multi-level decisions where participants may have weighed multiple factors (e.g., perceptual and/or semantic) to arrive at choice. This contrasts with the clear semantic boundaries that could be used to identify the odd-one-out in class-boundary triplets. Consistent with the RT results, we also found higher levels of cross-participant agreement (see Fig. 5C) in the class boundary condition compared to both the coarse-grained semantic and the fine-grained semantic conditions.
We used participants’ RTs as a proxy of their decision uncertainty. The human RTs for the individual triplets correlated positively with the level of cross-participant disagreement on those triplets (see Supplementary Material Sec. A.3 and Fig. 5B), in each of the abstraction settings (coarse-grained semantic: $r=0.325$, $p<0.001$; fine-grained semantic: $r=0.364$, $p<0.001$; class-boundary: $r=0.443$, $p<0.001$), suggesting that human RTs provided a reasonable approximation of decision uncertainty for model-to-human comparison.
Human-to-human alignment. We computed the human noise ceiling for each abstraction setting in Levels using a Leave-One-Out (LOO) cross-validation approach. In LOO, the agreement level for a triplet is computed as the average match-rate between a held-out participant’s response and the majority response of the remaining population. Thus, for a triplet with five responses, one response is held out and the remaining four comprise the population. The human-to-human reliability score is then calculated as the average agreement level across all triplets in the dataset.

### v4

@@ Abstract
Deep neural networks have achieved success across a wide range of applications, including as models of human behavior and neural representations in vision tasks [35, 64]. However, neural network training and human learning differ in fundamental ways, and neural networks often fail to generalize as robustly as humans do [45, 22], raising questions regarding the similarity of their underlying representations. What is missing for modern learning systems to exhibit more human-aligned behavior? We highlight a key misalignment between vision models and humans: whereas human conceptual knowledge is hierarchically organized from fine- to coarse-scale distinctions [e.g. 28], model representations do not accurately capture all these levels of abstraction. To address this misalignment, we first train a teacher model to imitate human judgments, then transfer human-aligned structure from its representations to refine the representations of pretrained state-of-the-art vision foundation models via finetuning. These human-aligned models more accurately approximate human behavior and uncertainty across a wide range of similarity tasks, including a new dataset of human judgments spanning multiple levels of semantic abstractions. They also perform better on a diverse set of machine learning tasks, increasing generalization and out-of-distribution robustness. Thus, infusing neural networks with additional human knowledge yields a best-of-both-worlds representation that is both more consistent with human cognitive judgments and more practically useful, thus paving the way toward more robust, interpretable, and human-aligned artificial intelligence systems.
@@
AI alignment, human semantic knowledge, representation learning, computer vision
@@ 1 Introduction
Figure 1: a: An example of the triplet odd-one-out task where a human and a neural network model choose the same (top) and a different (bottom) odd-one-out image respectively. b: The different parts of the AligNet framework depicted from end to end. First, we develop a teacher model of human judgments using the THINGS dataset (first panel). Second, we apply this model to ImageNet and cluster its latent representations into semantically meaningful categories (second panel). This allows us to generate arbitrarily many similarity judgments. Third, the obtained human-aligned similarity structure information is distilled into a student vision foundation model using a novel loss function (third panel). c: Representative human alignment (top panel) and machine learning downstream (bottom panel) results show significant performance improvements of the aligned over the non-aligned version of the ViT-B classifier (up to 123% of relative improvement for ML downstream transfer). d: 2D latent space projection for visualizing the change in the representations after alignment. While the representations of a standard ViT-B classifier model are unstructured and categories overlap, after alignment the representations are grouped into meaningful categories.
@@
While deep learning has recently driven rapid progress in areas of artificial intelligence such as natural language processing [47, 11] and computer vision [66, 17, 86, 39], even the best of these systems often fail in ways that humans would not [22, 46, 23, 79, 9, 56]. These failures have led to renewals [45, 9] of older arguments [19, 78] that neural networks lack essential ingredients of human intelligence. How can we build systems that produce more human-aligned behavior?
@@
Human perception is robust and generalizes across different visual settings [44, 70, 22]. However, model performance declines—often drastically—if the data distribution shifts between training and test sets [e.g., 31]. This lack of robustness in vision model representations poses a challenge for downstream applications that require generalization [e.g., 65]. In addition, humans tend to be well-calibrated—for example, when they are asked to judge visual similarity [70]—that is, humans’ (un)certainty tends to correlate with their (in)accuracy. AI systems, on the other hand, are often overconfident and show high certainty even when their predictions are incorrect [52]. Thus, many differences remain to be reconciled before we can ultimately achieve human-like artificial intelligence.
@@
Here, we highlight a key misalignment between humans and deep learning models that may underlie some of these differences: model representations tend to fail to capture the full multi-level conceptual structure of human knowledge. Although model representations successfully encode the local human-perceived similarity structure among closely-related entities that share (e.g. different dog breeds), the global relationships between concepts with more abstract semantic relations (e.g., dogs and fish, which are both animate but visually dissimilar) are modeled much less systematically. Human neural representations, however, are organized by global features like animacy [14, 28], and at multiple finer scales that capture nuanced semantic relationships [15, 14, 37]. Could the lack of global organization in the representations of deep learning models across levels of the conceptual hierarchy contribute to the aforementioned weaknesses of these models?
A challenge for addressing this misalignment is that collecting representative datasets of human judgments is challenging and expensive. We therefore propose a method for synthesizing simulated (approximately) human-aligned similarity judgments via a surrogate teacher model—a large foundation model which we align using an affine transformation [55] and uncertainty distillation on a small existing dataset [27]. We use this surrogate to produce the AligNet Dataset—human-aligned pseudolabels [cf. 48] from the surrogate model for triplets sampled from ImageNet [73] using a novel clustering-based data grouping method. We finetune various vision foundation models on AligNet using a novel similarity-space distillation objective. These models show substantially more human-aligned predictions on various cognitive science tasks—including Levels, a novel dataset of human semantic judgments reflecting multiple levels of abstraction. Furthermore, these aligned models show improved accuracy and out-of-distribution robustness across many downstream machine learning tasks, thus showing the improved generalizability of the aligned model representations. We release our aligned models and training and evaluation datasets.
@@
In summary, our work contributes to better understanding a key difference between artificial and natural intelligence. Moreover, our results illustrate a principle for aligning models to humans—focusing on the multi-scale relational structure of human knowledge—that may contribute to the more general problem of achieving human-aligned artificial intelligence.
@@ 2 Results
To build foundation models with more human-aligned behavior we inject additional supervision about human behavior into the model representations, using a surrogate teacher model: a vision foundation model whose representations are linearly transformed to approximate human judgments and uncertainty on the THINGS dataset [27]. We use clusters from this teacher model’s representations to sample triplets from ImageNet [73] and soft-label them using distances in the teachers representation space, then distill these soft-labels into the student representations via a Kullback-Leibler divergence objective. For details see Sec. 4.1.
2.1 Toward more human-aligned models
@@
THINGS Triplet Odd-one-out. We first validate that our teacher model performs well on the test data for the THINGS dataset [27] used in training; as expected, the teacher achieves high performance (61.7% accuracy, close to the human noise ceiling of 66.67%). We then align a variety of student models—trained with objectives ranging from image captioning to classification or self-supervised learning—using this teacher’s representations; all models show substantially improved human alignment on the THINGS tests (relative performance increases from 21.28 - 74.47%).
Other cognitive tasks. Our findings generalize across various object similarity tasks that are commonly used in the cognitive sciences—triplet odd-one-out task (relative performance increases up to $73.35\%$), Likert scale similarity ratings (up to $6.3$ fold increase in Spearman rank correlation coefficient), and multiple-arrangement tasks (up to $14.47$ fold increase in Spearman rank correlation coefficient). All performance increases are statistically significant at $\alpha=0.05$; for details see the SI.